### v1

# ZoFia: Zero-Shot Fake News Detection with Entity-Guided Retrieval and Multi-LLM Interaction

Abstract
The rapid spread of fake news threatens social stability and public trust, rendering its detection an imperative research priority. Although large language models (LLMs) excel at numerous natural language processing tasks with their remarkable contextual understanding and extensive prior knowledge, the time-bounded knowledge coverage and tendency for generating hallucination content reduce their reliability when handling fast-evolving news streams. Furthermore, models trained on existing static datasets also often lack the generalization needed for emerging news topics. To address these challenges, we propose ZoFia, a novel two-stage zero-shot fake news detection framework. First, we introduce Hierarchical Salience to quantify the importance of entities in the news content, and propose the SC-MMR algorithm to effectively select an informative and diverse set of keywords that serve as queries for retrieving up-to-date external evidence. Subsequently, a multi LLM interactive system, in which each agent assumes a distinct role, performs multi-view collaborative analysis and adversarial debate over the news text and its related information, and finally produces an interpretable and robust judgment. Comprehensive experiments on two public datasets demonstrate that ZoFia obviously outperforms existing zero-shot baselines and most of few-shot methods. Our codes will be open-sourced to facilitate related communities.
ZoFia: Zero-Shot Fake News Detection with Entity-Guided Retrieval and Multi-LLM Interaction
Lvhua Wu1,2thanks: Equal contribution., Xuefeng Jiang1,211footnotemark: 1, Sheng Sun1, Tian Wen1, Yuwei Wang1, Min Liu1,2thanks: Corresponding author. 1Institute of Computing Technology, Chinese Academy of Sciences 2University of Chinese Academy of Sciences {wulvhua24s,jiangxuefeng21b,liumin}@ict.ac.cn
1 Introduction
The rapid spread of fake news through social networks has become prevalent, posing severe threats to key domains such as politics Fisher et al. (2016), economy Bakir and McStay (2018), and livelihood Zhou and Zafarani (2020). The swift advancement of generative models further exacerbates this concern Chen and Shu (2023); Nan et al. (2024); Li et al. (2025, 2024). In this context, developing effective, efficient and interpretable fake news detection methods is imperative.
Fake news detection is formally defined as a binary classification problem. Prior to the rise of Large Language Models (LLMs), related research is predominantly led by training deep models via supervised learning Monti et al. (2019); Kaliyar et al. (2021). LLMs exhibit significant potential in information forensics due to their broad knowledge and strong contextual understanding acquired through pretraining on massive corpora Su et al. (2023). Current research approaches utilizing LLMs for detection generally fall into two categories. The first category employs LLMs as zero-shot reasoners, directly guiding models to judge and explain news authenticity through meticulously designed prompt engineering Jiang et al. (2022) or instruction tuning Pavlyshenko (2023). The second category treats LLMs as auxiliary tools, leveraging their powerful generative capabilities for information augmentation to enhance the performance and robustness of downstream compact classifiers Hu et al. (2024a).
However, directly applying LLMs to fake news detection poses significant challenges. The main problem is that LLMs have a knowledge cutoff Cheng et al. (2024), which means that an LLM’s internal knowledge becomes static after training, limiting its capacity to adapt to dynamic news streams. Furthermore, LLMs are prone to hallucinations Ji et al. (2023), frequently producing plausible but factually incorrect content, which directly undermines its reliability as a misinformation detector. More seriously, existing supervised methods generalize poorly to emerging topics Hoy and Koulouri (2022), and a single LLM used for zero-shot judgment often produces shallow reasoning, leading to superficial analysis.
To address these challenges systematically, we propose ZoFia, a training-free zero-shot framework for fake news detection. ZoFia decouples the complex detection task into two stages. First, it performs entity guided retrieval to integrate the LLM’s internal knowledge with real time external information and builds a comprehensive Multi-Source Information Matrix. Next, an multi-LLM interaction system conducts multidimensional analysis and adversarial debate over this matrix to deliver an explainable and robust judgment. The framework aims to compensate for the knowledge limits of LLMs while fully leveraging their inherent abilities of reasoning and interaction.
Our main contributions can be summarized as follows:
  • We propose ZoFia, a zero-shot fake news detection framework that combines entity-guided retrieval with multi-agent collaboration, effectively addressing LLMs’ knowledge cutoff and hallucinations.
  • We propose two novel algorithms, Hierarchical Salience and SC-MMR, to achieve accurate keyword extraction for news.
  • Experiments conducted on two public real-world datasets demonstrate that our proposed framework outperforms existing zero-shot models and most of few-shot models.
  • Code will be open-sourced to ensure the reproduction and facilitate related communities.
Figure 1: Overall architecture of our proposed ZoFia framework.
2 Related Work
2.1 LLM-based Fake News Detection
Using LLMs for fake news detection is a recent research trend, and current studies explores various capabilities of LLM. Hu et al. Hu et al. (2024a) position LLMs as auxiliary tools and propose that while LLMs cannot directly serve as detectors, they can provide multi-perspective guidance signals to smaller models. GenFEND Nan et al. (2024) focuses on information augmentation, which utilizes LLMs to generate diverse comments simulating different user groups to enrich related information. To address LLM knowledge limitations, Liu et al. Liu et al. (2024) combine internal sample retrieval with external real-time information retrieval, alleviating issues of vague and scarce information.
However, these methods seldom deeply explore other capabilities of LLMs, including inherent reasoning and role-playing. Distinct from prior studies, our framework guides LLMs to play different roles, collaboratively analyzing news from multiple expert perspectives and achieving better and more robust zero-shot detection.
2.2 Multi-agent Collaboration
Multi-agent collaboration as a new paradigm for enhancing LLM capabilities remains in its early stages. Its core concept involves integrating multiple agents to strengthen processing capacity for complex tasks. As pioneering work, Chateval Chan et al. (2023) demonstrated through agent team-based evaluation that this approach simultaneously improves robustness and accuracy in natural language generation. COLA Lan et al. (2024) applied this concept to zero-shot stance detection, assigning different roles to LLMs and combining a debate framework to enhance reasoning, yet failed to effectively mitigate hallucination problems. TruEDebate Liu et al. (2025) introduced multi-stage structured debate into fake news detection, but its implementation revealed a tendency toward premature convergence of viewpoints.
Our work uses a role-playing multi-agent paradigm, but also incorporates external knowledge retrieval and independent modular analysis to reduce hallucination and ensure diversity and independence of debate arguments.
3 Entity-Guided Retrieval
The efficiency of LLMs is constrained by the knowledge cutoff from training data Cheng et al. (2024). This limitation becomes especially apparent when processing rapidly changing real-time news streams, as misinformation usually propagates at rates and patterns that outpace the models’ training cycles Yi et al. (2025). Therefore, enhancing LLMs’ knowledge through external information retrieval is necessary.
We design the entity-guided retrieval stage to acquire real-time and reliable external knowledge for subsequent use by LLM experts. The framework consists of 4 sequential modules. The first 3 modules precisely extract and filter out a set of critical entities from the original news text as the keywords. These keywords are then used as queries for the final retrieval module, which retrieves from both Open Web and Wikipedia. This stage provides LLMs with the essential and up-to-date factual grounding and contextual information.
3.1 Entity Extractor
This module first uses a pre-trained BERT-NER Tjong Kim Sang and De Meulder (2003) model to perform named entity recognition (NER) on the news text. This process can be expressed as:
$\{(t_{i},e_{i},c_{i})\}_{i=1}^{N}=\mathcal{M}_{\text{BERT-NER}}(T),$ (1)
where $\mathcal{M}$ is the pre-trained model, $T$ is the input news text. $(t_{i},e_{i},c_{i})$ denotes the recognized entity triplet, $t_{i}$ is the entity token, $e_{i}$ is the entity label, and $c_{i}$ is the confidence score for the corresponding label, expressed as the conditional probability $c_{i}=P(e_{i}|t_{i},T;\mathcal{M}_{\text{BERT-NER}})$.
Due to the large number of recognized entities, we aggregate consecutive entity tokens to form new entity units $U_{e}$. Its confidence score $c(U_{e})$ is calculated by averaging the confidence scoresof all tokens in $U_{e}$, expressed as:
$c(U_{e})=\frac{1}{|U_{e}|}\sum_{t_{i}\in U_{e}}c_{i}.$ (2)
We establish a dynamic confidence threshold to select entities with high confidence scores. Starting from an initial confidence score $\lambda_{\text{init}}$, if the number of selected entities fails to meet the predetermined minimum $n_{\text{min}}$, the algorithm iteratively lowers the confidence score by $\Delta\lambda$ and repeats the selection, until at least $n_{\text{min}}$ entities are obtained.
3.2 Salience Scorer
This module aims to accurately quantify the importance of key entities in news texts, i.e. Entity Salience (ES) Dunietz and Gillick (2014). Although the prior study Bullough et al. (2024) has demonstrated that bi-encoder architectures can efficiently estimate entity salience, their performance is often limited by the semantic dilution Hou et al. (2021) problem. In lengthy or unevenly distributed texts, the core semantics of key entities are diluted by minor content, resulting in their true importance being severely underestimated.
We propose a novel Hierarchical Salience that avoids a single, coarse evaluation between an entity and the whole text. It decomposes an entity’s overall importance into two orthogonal and multiplicative components: Local Salience $\mathcal{S}_{\text{local}}$ and Global Salience $\mathcal{S}_{\text{global}}$. Local Salience measures the semantic alignment between the entity and its immediate context, and Global Salience measures how much the local context contributes to the text’s main content.
Formally, for a news text $T$ that is an ordered sequence of sentences, consider any entity $U_{i}$ that appears in sentence $s_{j}$. We define its local context as $\mathcal{C}(U_{i})=s_{j-1}\oplus s_{j}\oplus s_{j+1}$. Aligned with Bullough et al. (2024), we use a unified Sentence-BERT (SBERT) Reimers and Gurevych (2019) encoder $\mathcal{M}_{\text{SBERT}}(\cdot)$ to embed the entity $U_{i}$, its local context $\mathcal{C}(U_{i})$, and the full text $T$ to vectors $\mathbf{v}_{U_{i}}$, $\mathbf{v}_{\mathcal{C}(U_{i})}$, and $\mathbf{v}_{T}$. The hierarchical salience $\mathcal{S}_{\text{hier}}(U_{i})$ is derived from the product of these two components as
$\mathcal{S}_{\text{hier}}(U_{i})=\underbrace{\frac{\mathbf{v}_{U_{i}}\cdot\mathbf{v}_{\mathcal{C}(U_{i})}}{\|\mathbf{v}_{U_{i}}\|\cdot\|\mathbf{v}_{\mathcal{C}(U_{i})}\|}}_{\mathcal{S}_{\text{local}}(U_{i}\mid\mathcal{C}(U_{i}))}\cdot\underbrace{\frac{\mathbf{v}_{\mathcal{C}(U_{i})}\cdot\mathbf{v}_{T}}{\|\mathbf{v}_{\mathcal{C}(U_{i})}\|\cdot\|\mathbf{v}_{T}\|}}_{\mathcal{S}_{\text{global}}(\mathcal{C}(U_{i})\mid T)}.$ (3)
Hierarchical Salience provides a finer and more robust estimate of each entity’s importance to the news content. It serves as the key criterion for entity filtering in subsequent modules.
3.3 Keyword Selector
Input: $\mathcal{D}:\{U_{i}\to\mathcal{S}_{\text{Hier}}(U_{i})\}$: Mapping from entity to its salience score;
$\mathcal{M}_{\text{SBERT}}$: SBERT encoder for entity embeddings;
$\gamma$: Decay factor for termination criterion
Output: $\mathcal{U}_{\text{selected}}$: Final set of selected entities
$\mathcal{U}\leftarrow\text{keys}(\mathcal{D})$; $\mathbf{V}\leftarrow\{\mathcal{M}_{\text{SBERT}}(U_{i})\mid\forall U_{i}\in\mathcal{U}\}$
Compute pairwise similarity matrix $\mathbf{M}_{\text{sim}}$ from $\mathbf{V}$
if $\mathcal{U}_{\text{candidate}}=\emptyset$ then
 return $\mathcal{U}_{\text{selected}}$
$U^{*}\leftarrow\mathop{\arg\max}_{U_{i}\in\mathcal{U}_{\text{candidate}}}\mathcal{D}[U_{i}]$
$\mathcal{U}_{\text{selected}}\leftarrow\{U^{*}\}$; $\mathcal{U}_{\text{candidate}}\leftarrow\mathcal{U}\setminus\{U^{*}\}$
$\text{MMR}_{\text{prev}}^{*}\leftarrow 1.0$
while $\mathcal{U}_{\text{candidate}}\neq\emptyset$ do
 $k\leftarrow|\mathcal{U}_{\text{selected}}|$
 $\lambda_{k}\leftarrow\max(1.0\,,\,1.2-e^{0.2k-1.5}/e)$
 $\text{MMR}_{\text{best}}\leftarrow$-$\infty$; $U_{\text{best}}\leftarrow\text{null}$
 foreach $U_{i}\in\mathcal{U}_{\text{candidate}}$ do
    $\text{sim}_{\max}\leftarrow\max_{U_{j}\in\mathcal{U}_{\text{selected}}}\mathbf{M}_{\text{sim}}[i,j]$
    $\text{MMR}_{\text{curr}}\leftarrow\lambda_{k}\cdot\mathcal{D}[U_{i}]-(1-\lambda_{k})\cdot\text{sim}_{\max}$
    if $\text{MMR}_{\text{curr}}>\text{MMR}_{\text{best}}$ then
       $\text{MMR}_{\text{best}}\leftarrow\text{MMR}_{\text{curr}}$; $U_{\text{best}}\leftarrow U_{i}$
 if $\text{MMR}_{\text{best}}\leq\gamma\cdot\text{MMR}_{\text{prev}}^{*}$ then
    break
 $\text{MMR}_{\text{prev}}^{*}\leftarrow\text{MMR}_{\text{best}}$
 $\mathcal{U}_{\text{selected}}\leftarrow\mathcal{U}_{\text{selected}}\cup\{U_{\text{best}}\}$
 $\mathcal{U}_{\text{candidate}}\leftarrow\mathcal{U}_{\text{candidate}}\setminus\{U_{\text{best}}\}$
return $\mathcal{U}_{\text{selected}}$
Algorithm 1 Salience-Calibrated Maximal Marginal Relevance (SC-MMR)
This module aims to select an informative subset of keywords from the candidate entity set $\mathcal{U}_{\text{selected}}$ to mitigate the query drift Carpineto and Romano (2012) problem. However, this process faces two practical challenges. First, The hierarchical salience score $\mathcal{S}_{\text{hier}}(U_{i})$ is highly sensitive to context granularity, so a fixed screening threshold is not effective. Second, coreference in news introduces high–scoring entities with repeated semantics, which harms the diversity of the keyword set.
To optimize relevance and diversity under these constraints, We propose an improved MMR Carbonell and Goldstein (1998) algorithm, namely Salience-Calibrated MMR (SC-MMR) as shown in Algorithm 1. At the $k$-th iteration, SC-MMR evaluates the score of a candidate entity $U_{i}$ by
$\text{MMR}(U_{i})=\lambda_{k}\mathcal{S}_{\text{hier}}(U_{i})-(1-\lambda_{k})\max_{U_{j}\in\mathcal{U}_{\text{selected}}}\mathcal{S}(U_{i},U_{j}),$ (4a)
$\mathcal{S}(U_{i},U_{j})=\frac{\mathbf{v}_{U_{i}}\cdot\mathbf{v}_{U_{j}}}{\|\mathbf{v}_{U_{i}}\|\,\|\mathbf{v}_{U_{j}}\|}.$ (4b)
The core innovation of SC-MMR is a dynamic decaying weight $\lambda_{k}$. At the early stage, the weight places more emphasis on an entity’s own salience. As the number of selected entities $k$ increases, the weight adaptively decreases for high diversity and low redundancy. The computation of $\lambda_{k}$ is
$\lambda_{k}=\max\left[0.1\,,\,1.2-\frac{\exp(0.2k-0.5)}{e}\right].$ (5)
We further replace termination strategies that rely on a fixed score with a dynamic criterion based on the relative change of the MMR score. The iteration continues only when the next best candidate $\mathcal{U}_{k+1}^{*}$ satisfies $\text{MMR}(\mathcal{U}_{k+1}^{*})>\gamma\cdot\text{MMR}(\mathcal{U}_{k}^{*})$, where $\gamma$ is a decay factor. It lets the algorithm determine the stopping point according to score distribution and prevents a decline in the quality of entities caused by diminishing marginal utility.
3.4 Information Retrieval
This module uses the keyword set $\mathcal{K}=\{k_{1},k_{2},\ldots,k_{N}\}$ distilled in previous modules to build a comprehensive external knowledge base $\mathcal{E}$, serving as supplementary context for the subsequent detection stage. We implement a dual source retrieval that gathers information from the open web and Wikipedia in parallel, integrating the timeliness of the former with the authoritativeness of the latter.
Open Web. We compose all keywords into an aggregated query for the open web $Q_{\text{web}}$ with the following logic:
$Q_{\text{web}}=\left(\bigwedge_{i=1}^{N}k_{i}\right)\land\neg\texttt{news}\land\neg\texttt{wikipedia}.$ (6)
The operator $\land$ requires that results relate to all keywords, which enables strict cross-keyword verification. We explicitly exclude news and wikipedia sources. The former avoids retrieving the duplicate reports to prevent source contamination Deng et al. (2023), and the latter prevents redundancy with the subsequent Wikipedia retrieval. For each returned page, we only extract its summary and search snippet as the raw corpus.
Wikipedia. The summary section of a Wikipedia entry usually provides the most precise definition for an entity. However, a single keyword $k_{i}$ often corresponds to multiple Wikipedia entries, which introduces semantic ambiguity. We build a context-aware disambiguation mechanism that uses the original local context $\mathcal{C}(U_{i})$ of keyword $k_{i}$ in the news text to perform accurate matching.
When Wikipedia returns a list with $M$ candidate senses $\mathcal{O}(k_{i})=\{o_{1},o_{2},\ldots,o_{M}\}$, the mechanism examines each candidate $o_{m}$. It constructs a temporary modified context $\mathcal{C}(U_{i}\leftarrow o_{m})$ by replacing the original entity $U_{i}$ with the description text of $o_{m}$. The optimal sense $o_{i}^{*}$ of $U_{i}$ is defined as the option that maximizes the cosine similarity between the vector of the original context and the vector of the modified context:
$o_{i}^{*}=\mathop{\arg\max}_{o_{m}\in\mathcal{O}(k_{i})}\frac{\mathbf{v}_{\mathcal{C}(U_{i})}\cdot\mathbf{v}_{\mathcal{C}(U_{i}\leftarrow o_{m})}}{\|\mathbf{v}_{\mathcal{C}(U_{i})}\|\|\mathbf{v}_{\mathcal{C}(U_{i}\leftarrow o_{m})}\|}.$ (7)
where $\mathbf{v}$ denotes the context vectors embedded by the pretrained SBERT $\mathcal{M}_{\text{SBERT}}$. After identifying the unique entry, we extract the first 3 sentences of its summary as supplementary material.
4 Multi-LLM Interaction
Entity-guided retrieval supplies the framework with timely external information, which together with the prior knowledge of LLMs builds a Multi-Source Information Matrix. This stage employs a dual-state multi-LLM interaction system that fully exploits this matrix to mitigate hallucinations Feng et al. (2024) and ensure accurate detection.
This stage operates in two orthogonal interaction states. LLM Collaboration performs parallel analyses across multiple agents to reduce inferential variance. LLM Debate and Judgment introduces a adversarial debate to reduce systemic bias. These two states form a complete reasoning chain from mining divergent evidences to making a convergent judgment.
4.1 LLM Collaboration
LLM Collaboration state aims to reduce inferential variance. Through parallel analysis by multiple agents, it transforms Multi-Source Information Matrix from the previous stage into a structured evidence pool, which provides a stable decision basis for the subsequent adversarial debate. As illustrated in Figure 2, the matrix systematically integrates 4 information sources that are orthogonal and complementary in evidence distribution:
  • In-news Information: It comes from the original text of the target news and provides the basic core content.
  • Out-of-news Information: It comes from Open Web retrieval and offers the broadest instant facts.
  • LLM internal Knowledge: It comes from the model’s prior knowledge and provides generalized common sense.
  • LLM external Knowledge: It comes from Wikipedia retrieval and provides the most precise summary for the core entities.
The subsequent analysis decomposes the tasks and assigns them to agents with distinct roles, so that each agent focuses on a specific quadrant of the matrix for specialized processing.
Figure 2: The diagram of Multi-Source Information Matrix and the quadrants used by LLM agents.
4.1.1 Linguist
Following prior studies Shahid et al. (2022)Zhou and Zafarani (2018), the linguist agent is designed to systematically divide the text into 5 linguistic dimensions that are strongly associated with misinformation:
  • Sentence: Lexical complexity, sentence length, and formality of tone.
  • Word: Frequency of superlatives, affective language, and pronoun distribution.
  • Grammar: Patterns of reported speech, passive voice, and negation.
  • Emotion: Proportion of affective terms in the text and the headline, and the degree of incendiary tone.
  • Information Quality: Presence of clickbait, information overload, or context mismatch.
To maintain objective independence, each dimension is evaluated in an isolated session. For each dimension, LLM explicitly indicates whether it reflects the news is real or fake. This agent uses 2 quadrants of Multi-source Information Matrix: LLM internal knowledge and in-news information from the original news text.
4.1.2 Domain-Specific Expert
It operates in a dynamic and adaptive manner. The system first identifies the most relevant domain from the news text and assigns the agent a precise expert role, such as “economist” or “journalist”.
The expert with this role then analyses along the following 2 dimensions:
  • Knowledge Concordance: Examine all claims, viewpoints, and details for sound reasoning; identify departures from common sense.
  • Logical Integrity: Examine argument-to-conclusion coherence with domain-specific common sense; identify logical errors or unsupported leaps.
This agent uses 3 quadrants of Multi-source Information Matrix: in-news information, LLM internal knowledge, and LLM external knowledge from Wikipedia.
4.1.3 Claim Verification
Existing research Niu et al. (2024) demonstrates that claim-based fact-checking effectively serves as a reference for LLM-based detection. In our system, a serial pipeline composed of Claim Extractor and Claim Verifier deconstructs and verifies factual claims in the news text. These 2 agents use 3 quadrants of Multi-source Information Matrix: LLM internal knowledge, in-news information, and out-of-news information from Open Web.
Claim Extractor. $\mathcal{M}_{\text{extractor}}$ agent converts unstructured news text $T$ into a set of verifiable structured claims. Its function is formalized as follows:
$\mathcal{M}_{\text{extractor}}(T)\rightarrow\{q_{\text{core}},\{q_{\text{sub}_{1}},\dots,q_{\text{sub}_{m}}\}\}$ (8)
where $q_{\text{core}}$ is the core claim that determines the veracity of the news, and $q_{\text{sub}}$} is a collection of supporting subclaims. All outputs are restated as concise and objective declarative sentences.
Claim Verifier. This agent verifies each claim $q$ independently. To ensure precision and control cost Purwar et al. (2024), a simple retrieval augmented generation (RAG) Lewis et al. (2020) is implemented to build a highly relevant context $\mathcal{C}_{\text{rel}}(q)$ for each $q$ from the Open Web corpus $\mathcal{I}_{\text{web}}$.
The context consists of text chunks $c_{j}$ whose cosine similarity with the claim representation $\mathbf{v}_{q}$ exceeds a threshold $\theta_{\text{sim}}$:
$\mathcal{C}_{\text{rel}}(q)=\{c_{j}\in\text{top-k}(\mathcal{I}_{\text{web}},q)\mid\frac{\mathbf{v}_{q}\cdot\mathbf{v}_{c_{j}}}{\|\mathbf{v}_{q}\|\|\mathbf{v}_{c_{j}}\|}\geq\theta_{\text{sim}}\}.$ (9)
All decisions are strictly based on $\mathcal{C}_{\text{rel}}(q)$. The final output includes a clear label (“Supports”, “Refutes”, or “Not Enough Information”), a brief reasoning, and evidence directly quoted from $\mathcal{C}_{\text{rel}}(q)$ for suppressing hallucinations.
4.2 LLM Debate and Judgment
LLM Debate and Judgment state aims to suppress systemic bias. We introduce a multi-round adversarial debate framework that forces LLMs to explore both supporting and refuting views of the truthfulness of news equally. This design directly mitigates the thought degeneration (DoT) Liang et al. (2023) phenomenon that often appears in a single linear reasoning chain.
Input: $\mathbb{E}$: The complete evidence pool
Output: $D$: The final decision $\{\textit{Real},\textit{Fake}\}$
$\mathbb{H}\leftarrow\emptyset$; $A_{\text{con}}\leftarrow\text{null}$; $D\leftarrow\textit{Insufficient}$
while $D=\text{Insufficient}$ do
 $A_{\text{pro}}\leftarrow\mathcal{M}_{\text{pro}}.\text{GenerateArgument}(\mathbb{E},A_{\text{con}})$
 $A_{\text{con}}\leftarrow\mathcal{M}_{\text{con}}.\text{GenerateArgument}(\mathbb{E},A_{\text{pro}})$
 $\mathbb{H}\leftarrow\mathbb{H}\cup\{(A_{\text{pro}},A_{\text{con}})\}$
 $D\leftarrow\text{Judge.Assess}(\mathbb{H})$
return $D$
Algorithm 2 Dynamic Adversarial Debate
As shown in Algorithm 2, a pair of debate agents $\mathcal{M}_{\text{pro}}$ and $\mathcal{M}_{\text{con}}$ act as opposing reasoners based on the evidence pool $\mathbb{E}$. In each round, the active debater first rebuts the opponent’s last argument and then presents a new argument. After each exchange, a judge agent evaluates the debate history $\mathcal{H}$ and outputs a ternary judgment $D$.
This dynamic termination mechanism ensures that the debate stops once the information is sufficient for decision, effectively balancing the depth and efficiency of reasoning. The debate history $\mathbb{H}$ provides a traceable reasoning chain composed of pro. and con. arguments, making the final judgment highly interpretable.
5 Experiment
5.1 Experimental Setting
Herein we discuss main experimental settings.
Datasets. Following previous state-of-the-art work Liu et al. (2024), our experiments are conducted on two widely recognized fake news datasets: GossipCop and PolitiFact Shu et al. (2020). GossipCop focuses on entertainment, mainly Hollywood celebrity news; PolitiFact focuses on politics, drawing on fact-checks of U.S. political figures. Considering that some links in the original datasets have become invalid, we adopt subsets publicly released in Su et al. (2023).
Metrics and Baselines. We use the accuracy (Acc.) and F1-score (F1) as the main metrics. To reflect performance on detecting fake news, we compute the F1-Score for the fake class. ZoFia compares against two groups of baselines. The first group includes 5 advanced Few-Shot methods: MDFEND Nan et al. (2021), ARG Hu et al. (2024b), PSM Ni et al. (2020), DKFND Liu et al. (2024), and KPL Jiang et al. (2022), with metrics retrieved from the existing literature Jin et al. (2024); Hu et al. (2024c); Liu et al. (2024). The second group consists of 3 representative Zero-Shot methods that we implement for comparison: Standard Prompt, Auto-CoT Zhang et al. (2022), and ReAct Yao et al. (2023) equipped with a search API.
Implementation details. We set the dynamic threshold for entity extraction as $\lambda_{\text{init}}=0.8$ and $\Delta\lambda=0.1$ and the decay factor for MMR as $\gamma=0.5$. The maximum entries for Open Web retrieval is 10; for Wikipedia, we retrieve the first 3 sentences for each entry. The minimum similarity threshold for claim extraction is $\theta_{\text{sim}}=0.1$. The NER model is selected as dslim/bert-base-NER Tjong Kim Sang and De Meulder (2003), and the SBERT model is selected as all-MiniLM-L6-v2 Reimers and Gurevych (2019). All zero-shot methods we implement use DeepSeek-V3 DeepSeek‑AI (2024) with temperature set to 0.1 as the base model, and Brave API [^1] https://brave.com/search/api/ as the search API.
5.2 Main Experiment
The experimental results are presented in Table 1. ZoFia demonstrates exceptional performance on both datasets, consistently outperforming existing zero-shot and few-shot baselines. On the fact-intensive PolitiFact dataset, ZoFia’s advantages are particularly pronounced, with its accuracy (91.9%) and F1-score (89.8%) not only far exceeding other zero-shot methods but also surpassing all few-shot baselines by a substantial performance margin, proving its powerful efficacy in fact-checking tasks.
Table 1: Accuracy (Acc.) and F1-score (F1) comparison of few-shot / zero-shot methods on PolitiFact and GossipCop. The bold and underlined denote the best and second-best performance.
Method PolitiFact GossipCop Acc. F1 Acc. F1 MDFEND 65.5 62.3 41.2 40.2 ARG 74.0 67.1 61.4 42.3 PSM 70.0 49.1 77.4 41.7 DKFND 87.0 67.1 82.4 42.3 KPL 58.3 60.4 42.7 42.1 Standard Prompt 79.0 40.2 61.0 28.1 Auto-CoT 89.7 73.7 56.7 43.3 ReAct (Search API) 74.7 67.6 74.0 47.3 ZoFia (Ours) 91.9 89.8 78.5 61.2
The detection task on the GossipCop dataset presents greater challenges, as its content often proves difficult to verify due to subjectivity and factual ambiguity. In this scenario, ZoFia achieves the highest F1-score (61.2%), surpassing all baseline models. Although DKFND attains slightly higher accuracy, its F1-score is notably lower. In contrast, ZoFia’s superior F1-score demonstrates more balanced and robust detection performance across both true and fake news categories while maintaining high precision.
The experimental results clearly indicate that as a zero-shot method, ZoFia consistently outperforms the vast majority of few-shot and zero-shot models, showcasing its advancement and practical utility.
5.3 Ablation Study
To systematically assess the effectiveness of each component in ZoFia, we conduct extensive ablation on the GossipCop dataset.
Figure 3: Performance comparison of ZoFia’s keyword extraction with other extraction methods.
Effectiveness of Entity-Guided Retrieval. As shown in Figure 3, replacing our keyword extraction module with the classic TF-IDF Sparck Jones (1972) algorithm lowers the F1-Score by 9.1%. This finding emphasizes the need for our Hierarchical Salience. We also evaluate Sentence Salience Bullough et al. (2024) strategy, which retrieves a full sentence. The result shows an extreme imbalance: high Accuracy with low F1-Score. The reason is that using sentences as queries tends to retrieve pages that are highly similar to the original news. This source contamination Deng et al. (2023) induces confirmation bias in LLM agents.
Table 2: Results of ablation study on GossipCop
Component Acc. F1 ZoFia 78.5 61.2 w/o Open Web retrieval 71.7 58.1 w/o Wikipedia retrieval 77.7 59.9 w/o Open Web and Wikipedia retrieval 71.3 58.2 w/o linguist analysis 77.8 58.9 w/o expert analysis 76.8 57.3 w/o linguist and expert analyses 73.5 58.9 w/o claim verification 79.2 60.5 w/o debate 75.0 59.2
Component Analysis of Multi-LLM Interaction. As shown in Table 2, removing Open Web retrieval causes a marked decline, which demonstrates the importance of timely external knowledge. Performance decreases further when all external knowledge sources are removed. For LLM collaboration, removing the linguist or expert weakens the framework. The expert contributes more, with a 3.9% drop in F1-Score. Moreover, removing both causes cumulative degradation. Removing the debate reduces the F1-Score by 2.0%, demonstrating its role in suppressing single-view bias.
5.4 Sensitivity Analysis of SC-MMR
The dynamic weight $\lambda_{k}$ is the most important parameter of SC-MMR introduced in Section 3. We conduct a sensitivity analysis before designing it. First, $\lambda_{k}$ is fixed to 0.5, indicating equal emphasis on relevance and diversity in keyword selection. As shown in Figure 4, we observe that when the number of selected keywords $k\leq 6$, F1-Score remains stable at a relatively high level, but once $k>6$, F1-Score drops sharply. This pattern indicates that too many keywords introduce noise and hurt detection performance.
Figure 4: The effect of the number of keywords $k$ on performance (F1-Score).
Based on this finding, we designed $\lambda_{k}$ a decreasing function with an increasing decay rate, as shown in Eq. 5. It remains high for small $k$ to prioritize relevance, reaching approximately 0.45 at $k=6$, but rapidly decaying for $k>6$. This design lowers the MMR score and smoothly applies the relative termination rule. It allows SC-MMR to adapt the size of the keyword set while avoiding performance loss due to redundancy.
6 Conclusion
We propose ZoFia, a training-free zero-shot framework for fake news detection, to address the fundamental conflict between the static prior knowledge of large language models (LLMs) and the uncertainty of dynamic news streams. ZoFia uses a two-stage pipeline to tackle this challenge. In the first stage, our novel Hierarchical Salience and SC-MMR algorithms extract a set of informative and low-redundancy keywords from the news text, and then build a multi-source information matrix that integrates LLM internal and external knowledge with in-news and out-of-news information. In the second stage, a dual-state multi-LLM interaction processes this matrix. The agents first perform collaborative analysis to reduce inferential variance, and then conduct adversarial debate to suppress systemic bias, finally delivering an interpretable and robust judgment. Comprehensive experiments on two widely used datasets show that ZoFia outperforms existing zero-shot baselines and the majority of few-shot methods. Detailed ablation study further demonstrates the importance and effectiveness of each core component in the framework.
Limitations
Though ZoFia show strong detection capacities, its application and evaluation face multiple constraints. Due to copyright and privacy concerns, there has been a recent lack of high-quality, and continuously updated public datasets. It prevents us from evaluating ZoFia on the most recent news. Building the next generation of benchmark datasets that meet ethical standards and reflect real-world information dynamics is a critical step in this field.
From the framework’s perspective, ZoFia depends on multiple LLMs, which introduces nontrivial computation cost and inference latency. The efficiency of using external retrieval can also be improved. Since it is not our main focus, we integrate only a lightweight retrieval augmented generation (RAG) module in claim verification. Future work that adopts more advanced RAG architectures, such as re-ranking model and more advanced mechanism, may further strengthen the exploitation of external knowledge. We plan to conduct more comprehensive experimental benchmark evaluation over the detection capabilities of LLMs and Multi-agent system for fake news Guo et al. (2025); Kuntur et al. (2024); Jiang et al. (2024).
ZoFia currently focuses on the text modality. Modern misinformation increasingly appears in multi-modal form that combines images and text. Extending ZoFia to the multimodal domain has strong potential. One direction is to introduce vision language models (VLMs) as a dedicated visual expert. Another is to study how different modalities interact during the debate process to achieve effective fusion.